OpenStudy (anonymous):

plz help Suppose a cylindrical barrel falls off a ship and sinks to the bottom of the sea at a depth of 1.61 km. Assume that the seawater is incompressible, so that its density at the bottom of the sea is the same at the surface :1020 kg/m^3 If the pressure inside the barrel is 1.00 atm (it was sealed at sea level) and the total surface area of the barrel is 3.50 m^2 , find the NET INWARD FORCE acting on the surface of the barrel when it reaches the bottom of the ocean.

OpenStudy (praxer):

total pressure acting = atmospheric pressure + pressure with depth = 1.01*10^5 pa (atmospheric pressure) + h ro g ( pressure due to depth) =1.01*10^5 + 1610m*1020*10 =165.2*10^5pa =165.2atm now pressure = force /area force = pressure * area =165.2*10^5*3.50 =578.2*10^5 Newton.

OpenStudy (praxer):

@missylulu did you understand it.:)

OpenStudy (anonymous):

@praxer yeah thanks so much!!! :)))

OpenStudy (praxer):

$$\huge \frak \color{red}{WELCOME....\ddot\smile}$$